If C=1000+3/5y
S=4000+0.1y
G=2400+0.15y
X=2600x0.15y
M=1000+0.1y

1. determine the equilibrium income level of the given letter
2. What are the actual value at G, I GxM
1
Expert's answer
2019-08-29T09:36:40-0400

Y = C + G + I+Х-М

Where I = S

Y = 1000+3/5Y +2400 +0.15Y+4000+0.1Y+2600*0.15Y-1000-0.1Y

Y=3/5Y +2400 +0.15Y+4000+2600*0.15Y

Y = 0.75Y+6400+2600*0.15Y

Y-0.75Y-2600*0.15Y=6400

0.25Y-2600*0.15Y=6400

Y(0.25-2600*0.15)=6400

Y *(- 389.75)=6400

Y = -16.42

G=2400+0.15y =2400-2.46=2397.54

I=4000+0.1y = 3998.35

G*M=2397.54*(1000+0.1y)=2393603.23